# Survey on Neural Routing Solvers

### v1

@@ I Introduction
The vehicle routing problem (VRP) [25, 22] is a classic combinatorial optimization problem (COP) that seeks cost-minimizing routes for serving geographically distributed customers under specific constraints. Its scientific significance and broad practical impact have been demonstrated across various fields [128], such as transportation [24], logistics [45], and manufacturing [2]. As an NP-hard problem [70], VRPs cannot be solved to optimality in polynomial time, which has driven decades of research into heuristic algorithms to obtain high-quality approximations within acceptable computation time [69, 122]. However, designing effective heuristics requires substantial domain expertise and careful manual tuning, which poses significant challenges for real-world applications.
Efforts to automate heuristic design for combinatorial optimization have long been underway. A prominent direction is algorithm selection [113], which leverages features across problem instances to choose the most suitable algorithm for a given one. This idea has been extended to portfolio-based methods [51, 39], where a set of complementary algorithms is maintained and selectively applied to maximize performance for different instances [51, 39]. Another established paradigm is algorithm configuration, which aims to optimize algorithm performance for a target problem by automatically tuning parameters and combining modules [53, 41]. Despite their advances, these approaches remain confined to manually specified components and predefined parameter ranges, thus unable to discover or integrate novel algorithmic elements, which fundamentally limits potential performance gains.
@@
Category-Specific Insights: Transferring from heuristics to NRSs with corresponding categories.
@@
The growing literature on NRSs has been partially reviewed in several surveys, yet a significant gap remains. Surveys organized from a learning perspective [126, 127, 8, 66, 104, 13, 11, 114, 21, 123, 1, 95, 56] typically structure the field around DL techniques related to specific components, which cannot capture algorithmic structure and behavior of NRSs. Other surveys adopting a hybrid perspective [146, 6, 84, 136, 118, 71, 133, 153, 98, 134, 74, 156, 3] often introduce secondary attributes to define heuristic categories or restrict attention to NRSs with such attributes, leading to (1) incomplete coverage, forcing a residual “others” category, (2) ambiguous classification, resulting in self-contradictory taxonomies, or (3) limited scope, omitting NRSs without chosen attributes from discussion. Overall, these surveys lack a unified algorithm-level perspective on the field. A comparative summary is provided in Table I.
@@
Unified Algorithm-Level Review NRSs are inherently heuristic algorithms powered by DL models. Building on this understanding, a hierarchical taxonomy of NRSs is proposed from the perspective of heuristics, organized by how NRSs construct or improve solutions. This perspective clarifies the relationships among NRSs and highlights their progression from traditional heuristics. Furthermore, category-specific insights from heuristics are introduced to corresponding NRSs.
@@ II-E Proposed Taxonomy and Statistics
Figure 2: Hierarchical structure of the proposed NRS taxonomy. Each subcategory is presented with the proportion of existing studies. The statistics are obtained from Google Scholar between January 1, 2015, and November 17, 2025. The paper list is further filtered by content relevance and supplemented with relevant experience. Finally, there are a total of 439 papers, including 344 methods across various categories, as well as other related studies such as surveys and benchmarks. Note that an NRS may contribute to the counts of multiple subcategories, due to the adoption of multiple inference strategies.
@@ III-A1 Appending
MOTSP, MOCVRP Transformer RL P-MOCO [80] 2022 A multi-objective NRS
with a preference-conditioned model.
@@
CVRP, VRPTW MVMoE [154] 2024 An MoE-based NRS for multi-attribute VRPs.
OVRP, VRPB CaDA [72] 2025 A constraint-prompted dual-attention mechanism.
@@
NAR TSP GCN SL GCN [57] 2019 An NAR NRS with graph ConvNet.
AGNN RL DIMES [110] 2022 Proposing a differentiable parameterization
(Greedy) of the solution space.
@@ IV-A Small Neighborhood Methods
NAR AGNN RL DIMES [110] 2022 Proposing a differentiable parameterization of
(MCTS-k-opt) the solution space.
@@ IV-A2 Sequential Search
SDVRP one learned repair criteria.
Unrestricted CVRP, GAT, GRU RL EGATE [34] 2020 LNS with one destroy and one repair criteria
Direct LNS CVRPTW learned by a single model.
TSP, CVRP Transformer SL L2C-Insert [90] 2025 LNS with one handcrafted destroy and
(Iteration) one learned repair criteria.
@@
AR TSP, CVRP Transformer SL LEHD [88] 2023 Iterative random reconstructions of partial solutions
(RRC) via appending.
SIL [89] 2025 Iterative parallel reconstructions of partial solutions
(PRC) via appending and related iterative training without labels.
@@
TSP, CVRP, Transformer RL LCP [61] 2021 Iterative re-decompositions and revisions.
Restricted PCTSP
Direct LNS (A)TSP, OP,
@@
/ CVRP(TW), VRPMPD Transformer RL L2D [76] 2021 Iterative subproblem selection and optimization.
LSTM, RL RBG [157] 2022 Iterative re-partitioning, merging, and re-solving.
@@ IV-B1 Direct LNS
TABLE VI: Representative Improvement-based Population-based NRSs
Search Generation Solvable Backbone Learning Method Year Remarks
Space Paradigm VRPs Paradigm
Transformer RL COMPASS [14] 2023 Latent space search with CMA-ES.

### v2

@@ I Introduction
The vehicle routing problem (VRP) [25, 22] is a classic combinatorial optimization problem (COP) that seeks cost-minimizing routes for serving geographically distributed customers under specific constraints. Its scientific significance and broad practical impact have been demonstrated across various fields [129], such as transportation [24], logistics [44], and manufacturing [2]. As an NP-hard problem [71], VRPs cannot be solved to optimality in polynomial time, which has driven decades of research into heuristic algorithms to obtain high-quality approximations within acceptable computation time [70]. However, designing effective heuristics requires substantial domain expertise and careful manual tuning, which poses significant challenges for real-world applications.
Efforts to automate heuristic design for combinatorial optimization have long been underway. A prominent direction is algorithm selection [115], which leverages features across problem instances to choose the most suitable algorithm for a given one. This idea has been extended to portfolio-based methods [52, 39], where a set of complementary algorithms is maintained and selectively applied to maximize performance for different instances [52, 39]. Another established paradigm is algorithm configuration, which aims to optimize algorithm performance for a target problem by automatically tuning parameters and combining modules [54]. Despite their advances, these approaches remain confined to manually specified components and predefined parameter ranges, thus unable to discover or integrate novel algorithmic elements, which fundamentally limits potential performance gains.
@@
Category-Specific and Category-agnostic Insights: Transferring from heuristics to NRSs.
@@
The growing literature on NRSs has been partially reviewed in several surveys, yet a significant gap remains. Surveys organized from a learning perspective [127, 128, 8, 67, 106, 13, 11, 116, 21, 124, 1, 96, 57] typically structure the field around DL techniques related to specific components, which cannot capture algorithmic structure and behavior of NRSs. Other surveys adopting a hybrid perspective [147, 6, 85, 137, 120, 72, 134, 155, 99, 135, 75, 158, 3] often introduce secondary attributes to define heuristic categories or restrict attention to NRSs with such attributes, leading to (1) incomplete coverage, forcing a residual “others” category, (2) ambiguous classification, resulting in self-contradictory taxonomies, or (3) limited scope, omitting NRSs without chosen attributes from discussion. Overall, these surveys lack a unified algorithm-level perspective on the field. This makes it difficult to clearly characterize algorithmic commonalities and differences among NRSs, and to provide systematic insights and future directions for each category and the field. A comparative summary is provided in Table I.
@@
Unified Algorithm-Level Review NRSs are inherently heuristic algorithms powered by DL models. Building on this understanding, a hierarchical taxonomy of NRSs is proposed from the perspective of heuristics, organized by how NRSs construct or improve solutions. This perspective clarifies the relationships among NRSs and highlights their progression from traditional heuristics. Furthermore, both category-specific and category-agnostic insights from heuristics are introduced into NRSs to guide further development.
@@ II-E Proposed Taxonomy and Statistics
Figure 2: Hierarchical structure of the proposed NRS taxonomy. Each subcategory is presented with the proportion of existing studies. The statistics are obtained from Google Scholar between January 1, 2015, and November 17, 2025. The paper list is further filtered by content relevance and supplemented with relevant experience and subsequent publications. Finally, there are 439 papers, including 344 methods across categories, as well as other related studies such as surveys and benchmarks. Note that an NRS may be counted in multiple subcategories due to its use of multiple inference strategies.
@@ III-A1 Appending
MOTSP, Transformer RL P-MOCO [81] 2022 A multi-objective NRS
MOCVRP with a preference-conditioned model.
@@
TSP, CVRP(TW) Transformer RL PolyNet [47] 2025 Complementary solution strategies for diversity.
@@
CVRP, VRPTW, MVMoE [156] 2024 An MoE-based NRS for multi-attribute VRPs.
OVRP, VRPB, CaDA [73] 2025 A constraint-prompted dual-attention mechanism.
@@
(Variants of)
(A)CVRP(TW),
OVRP, PDVRP, Transformer RL URS [154] 2025 An NRS capable of solving more than 100 VRP
VRPB(P), VRPL, variants with a single model.
MDVRP, PCVRP
@@
TSP GCN SL GCN [58] 2019 An NAR NRS with graph ConvNet.
NAR AGNN RL DIMES [112] 2022 Differentiable parameterizations of solution spaces.
@@ IV-A Small Neighborhood Methods
NAR AGNN RL DIMES [112] 2022 Differentiable parameterizations of solution spaces.
@@
  • $\ddagger$
    The NRS supports multiple inference strategies and currently employs an improvement-based MCTS-k-opt inference.
@@ IV-A2 Sequential Search
Unrestricted SDVRP one learned repair criteria.
Direct LNS CVRP(TW) GAT, GRU RL EGATE [34] 2020 LNS with a learned criteria for both destroy and repair.
@@
TSP, CVRP Transformer SL L2C-Insert [91] 2025 LNS with a handcrafted destroy and a learned repair criteria.
CVRP(TW), Transformer RL NDS [49] 2025 LNS with a learned destroy and a handcrafted repair criteria.
PCVRP
@@
/ CVRP(TW) GAT, GRU SL L2Seg [103] 2026 A decomposition technique to accelerate iterative solvers.
Transformer
Restricted AR TSP, CVRP Transformer SL LEHD [89] 2023 Random reconstructions of partial solutions via appending.
Direct LNS SIL [90] 2025 Parallel reconstructions of partial solutions via appending.
@@
(PC)TSP, CVRP Transformer RL LCP [62] 2021 Iterative re-decompositions and revisions.
(A)TSP, OP,
@@
/ CVRP(TW), Transformer RL L2D [77] 2021 Iterative subproblem selection and optimization.
VRPMPD LSTM, RL RBG [159] 2022 Iterative re-partitioning, merging, and re-solving.
@@ V Population-Based Methods
TABLE VI: Representative Improvement-based Population-based NRSs
Search Generation Solvable Backbone Learning Method Year Remarks
Space Paradigm VRPs Paradigm
Transformer RL COMPASS [14] 2023 Latent space search with CMA-ES.
@@ VIII Conclusions
Acknowledgments
The authors would like to thank K. Li, Z. Xie, F. Luo, and J. Hou for valuable support in the experimental studies.